OpenStudy (anonymous):

4x+1=5+2(2x-4) A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

there is an infinate amount of solutions so your answer is b

OpenStudy (anonymous):

thxs,i got more questions.only 2 more,ill give medals.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

a+5=1over5(15+5a) A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

a i wont give it though but as long as u dont have to solve it then it is A

OpenStudy (anonymous):

thank you so much, one more and thts it

OpenStudy (anonymous):

kk ask away

OpenStudy (anonymous):

Which expressions would complete this equation so that it has infinitely many solutions? 8 + 2(8x – 6) = Choose exactly two answers that are correct. A. 2(4x + 7) B. 9x – 10 C. 16x – 4 D. 4(4x – 1)

OpenStudy (anonymous):

oh crap...it came out wrong

OpenStudy (anonymous):

8 + 2(8x – 6) = Choose exactly two anwser that are correct. A.2(4x+7) B.9X-10 C.16X-4 D.4(4X-1)

OpenStudy (anonymous):

8+2(8X-6)=*

OpenStudy (anonymous):

A and D

OpenStudy (anonymous):

hope this helped :)